# Spurious-Aware Prototype Refinement for Reliable Out-of-Distribution Detection

### v1

@@ 6 Discussion, Limitations, and Broader Impact
Experimental results consistently show SPROD’s superior performance across five diverse SP-OOD benchmarks, including our newly introduced Animals MetaCoCo dataset. Our evaluations also reveal that fine-tuning the feature backbone on in-distribution data can degrade SP-OOD detection performance. Furthermore, investigations into scoring mechanisms highlighted the advantages of distance-based approaches over softmax-based scoring for SP-OOD detection, particularly affirming the design choices in SPROD.
SPROD offers a scalable solution for improving robustness in OOD detection. While demonstrating significant advancements, SPROD’s current formulation relies on class labels from the in-distribution training data to construct class-conditional prototypes. This reliance aligns with the typical assumptions of the spurious correlation setting, which presumes access to data samples with spuriously correlated labels. Future work could explore theoretical justifications for the observed benefits of its generative-like scoring mechanism or investigate more expressive methods for modeling class-conditional distributions.
@@ Acknowledgments
Appendix A Spurious Correlation in OOD Detection: Literature Review
@@
Outlier exposure based methods include RONF [21], which improves synthetic outlier generation and model fine-tuning using only in-distribution (ID) data. It introduces Boundary Feature Mixup to create more realistic virtual outliers by interpolating near decision boundaries, and Optimal Parameter Learning to suppress spurious feature learning during training. At inference, it uses a custom Energy with Energy Discrepancy score to better separate ID from OOD samples without relying on external OOD datasets. Similarly, KIRBY [22] generates hard negative samples by removing class-discriminative regions identified via Class Activation Maps (CAM) and inpainting these regions with background-like content, creating semantically degraded but visually plausible images as surrogate OOD examples. A lightweight rejection network trained on features from both clean and modified images enables strong OOD detection without requiring real OOD data or backbone retraining. Although not directly targeting spurious correlations, ImOOD [23] evaluates robustness under spurious settings, focusing on long-tailed datasets where class imbalance biases OOD detection toward frequent (head) classes. It learns a bias correction term to shift OOD scores per input, improving separation especially for rare (tail) classes.
@@
Post-hoc methods operate after model training to improve OOD detection. Projection Regret (PR) [27] uses partial diffusion and denoising to project inputs onto the in-distribution manifold, measuring semantic novelty by comparing the original image with its projection. A second projection step recursively removes background bias, isolating true semantic differences and enabling object-level OOD detection focused on meaningful changes rather than background similarity. CoVer [28] enhances detection by evaluating model confidence across multiple corrupted versions of the same input (e.g., fog, blur, contrast shifts). Instead of relying on a single prediction, it averages confidence scores over these variants, improving robustness.
@@
A.1 Comparison with Related Methods
Most existing OOD detection methods are not directly comparable to SPROD, as they often rely on different assumptions, such as access to auxiliary OOD data, or necessitate retraining the feature backbone with specific objectives. These design choices contrast with SPROD’s post-hoc nature and its objectives of simplicity, efficiency, and broad applicability without model retraining. Nonetheless, to provide a broader context, we include results from a selection of methods that, while not perfectly aligned, have been evaluated under similar spurious OOD conditions. For fairness and consistency, we report the results as published in their original papers, rather than re-implementing them, thereby representing each method according to its best-reported performance. As shown in Table 3, under the most comparable conditions presented, SPROD achieves highly competitive performance, notably without requiring model retraining (when using a ResNet-18 backbone, as applicable to some comparisons) or access to an auxiliary text modality (in the case of methods leveraging CLIP).
@@ A.1 Comparison with Related Methods
Appendix B Dataset Details and Examples
Datasets available for studying spurious correlations are generally limited. In the context of OOD detection, this limitation becomes even more pronounced, as the task requires datasets with spurious OOD samples. As a result, the datasets employed must be thoughtfully considered to ensure meaningful evaluation. In this study, our goal is to address a broader range of spurious features present in data, beyond just background features. A distinctive aspect of our experimental design is the inclusion of three additional settings to explore:
@@ Appendix B Dataset Details and Examples
[…]
@@
  • AnimalsMetaCoCo: To create a multi-class, multi-valued spurious attribute setting that reflects realistic and challenging scenarios, we selected a subset of 26 animal classes from the MetaCoCo dataset [58]. The samples were first cleaned through relabeling, removal of duplicates, and deletion of irrelevant images. Subsequently, we defined subattributes to create a total of 8 major concepts for modeling attribute imbalance and spurious features, as illustrated in the first Figure of the paper. We introduce this new dataset as AnimalsMetaCoCo, a refined subset tailored for multi-class, multi-valued spurious feature scenarios. Details of the classes, attributes, and group sample distributions are provided in Table 7. This dataset represents a more realistic scenario, where each class is associated with at least one spurious feature similar to those present in ID data. While the strength of spurious correlations may not be as pronounced as in the other three datasets, AnimalsMetaCoCo introduces several unique and challenging aspects. Specifically, its spurious attributes can take on multiple distinct values, each exhibiting varying degrees of imbalance across classes. Moreover, the class distributions themselves are inherently imbalanced. To construct Spurious OOD scenarios, we adopt a leave-2-out strategy: in each round, two classes are treated as Spurious OOD, characterized by semantic shifts while retaining at least one spurious attribute shared with ID data, and the remaining classes serve as ID. This setup significantly increases the difficulty of the detection task due to overlapping spurious patterns across environments.
@@
Appendix C Performance on Non-Spurious OOD Datasets

### v2

@@ 6 Discussion
Experimental results consistently demonstrate the superiority of SPROD across ten convolutional and transformer-based backbones and five diverse SP-OOD benchmarks, including the newly introduced Animals MetaCoCo dataset. Our evaluations also reveal that fine-tuning the feature backbone on ID data can degrade SP-OOD detection performance. Furthermore, investigations into scoring mechanisms highlighted the advantages of distance-based approaches over softmax-based scoring for SP-OOD detection, particularly affirming the design choices in SPROD.
SPROD offers a scalable solution for improving robustness in OOD detection. While demonstrating significant advancements, SPROD’s current formulation relies on class labels from the ID training data to construct class-conditional prototypes. This reliance aligns with the typical assumptions of the spurious correlation setting, which presumes access to data samples with spuriously correlated labels. SPROD is not always the single best-performing method on every metric or in every setting (particularly in the conventional setting), but it remains competitive with the best-performing approaches. Such trade-offs are common in robustness research, where minor reductions in overall accuracy are often accepted to attain higher worst-group performance. Future work could explore theoretical justifications for the robustness of the generative-like scoring mechanism or investigate more expressive approaches for modeling class-conditional distributions.
@@ Acknowledgments
Appendix A Theoretical Analysis of SPROD
SPROD employs a two-step prototype refinement strategy to approximate group-specific representations within each class. The first step creates new prototypes for misclassified training samples, motivated by the observation that minority group instances (those that deviate from dominant spurious patterns) are more prone to misclassification. However, some correctly classified minority samples may still be incorrectly assigned to biased prototypes.
To address this, the second step reassigns each training sample to its nearest prototype and recalculates prototypes based on these updated assignments. This reassignment helps produce cleaner subgroup representations by reducing the influence of spurious-majority samples on minority feature structure.
By applying these two steps, SPROD aims to obtain nearly pure group-specific prototypes. In the remainder of this section, we provide a theoretical formulation to explain why such subgroup-specific prototypes improve robustness to spurious OOD (SP-OOD) samples.
A.0.1 Feature Decomposition
Without loss of generality, we assume a binary classification problem with classes $c\in\{0,1\}$. Each input sample $x_{i}$ is mapped to a normalized feature embedding $z_{i}\in\mathbb{R}^{d}$. We consider a general decomposition of $z_{i}$ into three semantically distinct components: spurious, core, and irrelevant features. We model $z_{i}$ as a linear combination of three functionally distinct components that span the embedding space:
$z_{i}=\sum_{j=1}^{n_{u_{c}^{\text{sp}}}}\alpha_{i,j}^{\text{sp}}\,\vec{\mathbf{u}}_{c,j}^{\text{sp}}+\sum_{j=1}^{n_{u_{c}^{\text{core}}}}\beta_{i,j}^{\text{core}}\,\vec{\mathbf{u}}_{c,j}^{\text{core}}+\sum_{j=1}^{n_{u^{\text{irr}}}}\gamma_{i,j}^{\text{irr}}\,\vec{\mathbf{u}}_{j}^{\text{irr}},$
where the sets $\{\vec{\mathbf{u}}_{c,j}^{\text{sp}}\}$, $\{\vec{\mathbf{u}}_{c,j}^{\text{core}}\}$, and $\{\vec{\mathbf{u}}_{j}^{\text{irr}}\}$ form orthonormal bases for the spurious, core, and irrelevant subspaces, respectively. Each set of coefficients $\{\alpha_{i,j}^{\text{sp}}\}$, $\{\beta_{i,j}^{\text{core}}\}$, and $\{\gamma_{i,j}^{\text{irr}}\}$ is specific to instance $i$.
These subspaces span mutually exclusive semantic roles:
  • Spurious subspace: captures features correlated with the label during training but not causally predictive.
  • Core subspace: captures features that are causally predictive of the label.
  • Irrelevant subspace: captures features unrelated to the task.
By definition, the basis vectors across subspaces are orthogonal: if an irrelevant basis vector were correlated with a core basis, it would be predictive and thus belong in the core subspace.
This decomposition can be expressed compactly in matrix form:
$z_{i}=U_{c}^{\text{sp}}\,\vec{\boldsymbol{\alpha}}_{i}^{\text{sp}}+U_{c}^{\text{core}}\,\vec{\boldsymbol{\beta}}_{i}^{\text{core}}+U^{\text{irr}}\,\vec{\boldsymbol{\gamma}}_{i}^{\text{irr}},$
where each $U^{(\cdot)}\in\mathbb{R}^{d\times n_{u}^{(\cdot)}}$ is a matrix of orthonormal basis vectors for the corresponding subspace, and each coefficient vector (e.g., $\vec{\boldsymbol{\alpha}}_{i}^{\text{sp}}\in\mathbb{R}^{n_{u_{c}^{\text{sp}}}}$) represents the coordinates of $z_{i}$ in that subspace.
A.0.2 Distributional Assumptions
We assume the coefficient vectors for each instance $z_{i}$ are drawn from class-conditional distributions:
$\vec{\boldsymbol{\alpha}}_{i}^{\text{sp}}\sim P_{c}^{\text{sp}},\quad\vec{\boldsymbol{\beta}}_{i}^{\text{core}}\sim P_{c}^{\text{core}},$
where $P_{c}^{\text{sp}}$ and $P_{c}^{\text{core}}$ are distributions over the spurious and core subspace coefficients for class $c$.
The irrelevant component is shared across classes:
$\vec{\boldsymbol{\gamma}}_{i}^{\text{irr}}\sim P^{\text{irr}}.$
A.0.3 Group Definitions
Within each class $c$, we define majority and minority groups based on the alignment of spurious features.
Majority group samples $z_{i}\in\mathcal{S}_{c}^{\text{maj}}$ satisfy:
$\vec{\boldsymbol{\alpha}}_{i}^{\text{sp}}\sim P_{c}^{\text{sp}},\quad\vec{\boldsymbol{\beta}}_{i}^{\text{core}}\sim P_{c}^{\text{core}}.$
Minority group samples $z_{i}\in\mathcal{S}_{c}^{\text{min}}$ satisfy:
$\vec{\boldsymbol{\alpha}}_{i}^{\text{sp}}\sim P_{1-c}^{\text{sp}},\quad\vec{\boldsymbol{\beta}}_{i}^{\text{core}}\sim P_{c}^{\text{core}}.$
Let $C_{c}$ denote the number of samples in class $c$, and $|\mathcal{S}_{c}^{\text{maj}}|$ the number of majority samples. Then:
$r_{c}=\frac{|\mathcal{S}_{c}^{\text{maj}}|}{C_{c}}\in[0,1]$
quantifies the class-conditional spurious correlation strength.
A.0.4 OOD Sample Definition
OOD samples may contain components beyond the core, spurious, and irrelevant subspaces. To model this, we extend the feature decomposition to include an additional subspace $U^{\text{ext}}$, capturing external factors not observed during training. We define an OOD sample as:
$z_{\text{OOD}}=U_{c}^{\text{sp}}\vec{\boldsymbol{\alpha}}_{\text{OOD}}^{\text{sp}}+U_{c}^{\text{irr}}\vec{\boldsymbol{\gamma}}_{\text{OOD}}^{\text{irr}}+U^{\text{ext}}\vec{\boldsymbol{\delta}}_{\text{OOD}}^{\text{ext}},$
where $\vec{\boldsymbol{\alpha}}_{\text{OOD}}^{\text{sp}}\sim P_{c}^{\text{sp}}$, $\vec{\boldsymbol{\gamma}}_{\text{OOD}}^{\text{irr}}\sim P^{\text{irr}}$, and $\vec{\boldsymbol{\delta}}_{\text{OOD}}^{\text{ext}}$ is unconstrained. Hard OOD samples may also include core components drawn from a shifted distribution $\vec{\boldsymbol{\beta}}_{\text{OOD}}^{\text{core}}\sim Q^{\text{core}}\neq P^{\text{core}}_{c}$. Although it is possible to analyze the general form with core and external components, for simplicity we focus on near-OOD samples with no core or external components, i.e.,
$z_{\text{OOD}}=U_{c}^{\text{sp}}\vec{\boldsymbol{\alpha}}_{\text{OOD}}^{\text{sp}}+U_{c}^{\text{irr}}\vec{\boldsymbol{\gamma}}_{\text{OOD}}^{\text{irr}},\quad\text{with }\vec{\boldsymbol{\beta}}_{\text{OOD}}^{\text{core}}=\vec{0},\;\vec{\boldsymbol{\delta}}_{\text{OOD}}^{\text{ext}}=\vec{0}.$
We define spurious-OOD groups based on the source of the spurious component:
$\mathcal{S}_{c}^{\text{OOD}}=\left\{z_{\text{OOD}}\,\middle|\,\vec{\boldsymbol{\alpha}}_{\text{OOD}}^{\text{sp}}\sim P_{c}^{\text{sp}},\;\vec{\boldsymbol{\gamma}}_{\text{OOD}}^{\text{irr}}\sim P^{\text{irr}}\right\},\quad c\in\{0,1\}.$
These near-OOD samples resemble class $c$ only through spurious features and match the in-distribution (ID) irrelevant component distribution.
A.0.5 Prototype Calculation
We define the expected coefficient vectors for each subspace:
$\vec{\boldsymbol{\mu}}_{c}^{\text{core}}=\mathbb{E}_{\vec{\boldsymbol{\beta}}\sim P_{c}^{\text{core}}}[\vec{\boldsymbol{\beta}}],\quad\vec{\boldsymbol{\mu}}_{c}^{\text{sp}}=\mathbb{E}_{\vec{\boldsymbol{\alpha}}\sim P_{c}^{\text{sp}}}[\vec{\boldsymbol{\alpha}}],\quad\vec{\boldsymbol{\mu}}^{\text{irr}}=\mathbb{E}_{\vec{\boldsymbol{\gamma}}\sim P^{\text{irr}}}[\vec{\boldsymbol{\gamma}}].$
Using these, we define the majority and minority subgroup prototypes for class $c\in\{0,1\}$ as:
$\vec{\mathbf{p}}_{c}^{\text{maj}}=U_{c}^{\text{sp}}\vec{\boldsymbol{\mu}}_{c}^{\text{sp}}+U_{c}^{\text{core}}\vec{\boldsymbol{\mu}}_{c}^{\text{core}}+U^{\text{irr}}\vec{\boldsymbol{\mu}}^{\text{irr}},$
$\vec{\mathbf{p}}_{c}^{\text{min}}=U_{c}^{\text{sp}}\vec{\boldsymbol{\mu}}_{1-c}^{\text{sp}}+U_{c}^{\text{core}}\vec{\boldsymbol{\mu}}_{c}^{\text{core}}+U^{\text{irr}}\vec{\boldsymbol{\mu}}^{\text{irr}}.$
The overall class prototype is a convex combination of subgroup prototypes:
$\vec{\mathbf{p}}_{c}=r_{c}\vec{\mathbf{p}}_{c}^{\text{maj}}+(1-r_{c})\vec{\mathbf{p}}_{c}^{\text{min}},$
where $r_{c}\in[0,1]$ denotes the proportion of majority group samples in class $c$.
A.0.6 Bias in Prototype Distances Under Strong Spurious Correlation
In standard prototype-based OOD detection, the class prototype $\vec{\mathbf{p}}_{c}$ is a convex combination of majority and minority subgroup prototypes. In spurious-dominated regimes where $r_{c}\approx 1$, we have:
$\vec{\mathbf{p}}_{c}\approx\vec{\mathbf{p}}_{c}^{\text{maj}}=U_{c}^{\text{sp}}\vec{\boldsymbol{\mu}}_{c}^{\text{sp}}+U_{c}^{\text{core}}\vec{\boldsymbol{\mu}}_{c}^{\text{core}}+U^{\text{irr}}\vec{\boldsymbol{\mu}}^{\text{irr}}.$
Case 1: Spurious-OOD Sample. Let $z_{\text{OOD}}\in\mathcal{S}_{c}^{\text{OOD}}$ be an OOD sample with spurious alignment to class $c$, but no core features:
$z_{\text{OOD}}=U_{c}^{\text{sp}}\vec{\boldsymbol{\alpha}}_{\text{OOD}}+U^{\text{irr}}\vec{\boldsymbol{\gamma}}_{\text{OOD}},\quad\text{with }\vec{\boldsymbol{\alpha}}_{\text{OOD}}\sim P_{c}^{\text{sp}},\;\vec{\boldsymbol{\gamma}}_{\text{OOD}}\sim P^{\text{irr}}.$
The expected squared distance to the biased class prototype becomes:
$\mathbb{E}\left[\|z_{\text{OOD}}-\vec{\mathbf{p}}_{c}\|^{2}\right]=\underbrace{\|\vec{\boldsymbol{\mu}}_{c}^{\text{core}}\|^{2}}_{\text{core bias}}+\underbrace{\operatorname{Tr}(\Sigma_{c}^{\text{sp}})}_{\text{spurious variance}}+\underbrace{\operatorname{Tr}(\Sigma^{\text{irr}})}_{\text{irrelevant variance}}.$
Case 2: Minority In-Distribution Sample. Now consider an ID sample from the minority group:
$z_{\text{min}}=U_{c}^{\text{sp}}\vec{\boldsymbol{\alpha}}_{\text{min}}+U_{c}^{\text{core}}\vec{\boldsymbol{\beta}}_{\text{min}}+U^{\text{irr}}\vec{\boldsymbol{\gamma}}_{\text{min}},$
with:
$\vec{\boldsymbol{\alpha}}_{\text{min}}\sim P_{1-c}^{\text{sp}},\quad\vec{\boldsymbol{\beta}}_{\text{min}}\sim P_{c}^{\text{core}},\quad\vec{\boldsymbol{\gamma}}_{\text{min}}\sim P^{\text{irr}}.$
Its expected squared distance to the majority prototype is:
$\mathbb{E}\left[\|z_{\text{min}}-\vec{\mathbf{p}}_{c}^{\text{maj}}\|^{2}\right]=\underbrace{\|\vec{\boldsymbol{\mu}}_{1-c}^{\text{sp}}-\vec{\boldsymbol{\mu}}_{c}^{\text{sp}}\|^{2}}_{\text{spurious bias}}+\underbrace{\operatorname{Tr}(\Sigma_{c}^{\text{core}})}_{\text{core variance}}+\underbrace{\operatorname{Tr}(\Sigma^{\text{irr}})}_{\text{irrelevant variance}}.$
Key Insight: Although $z_{\text{min}}$ is an ID sample, its distance to the biased prototype includes a spurious bias term, while the OOD sample differs only in the core direction. Depending on the relative magnitudes of the core and spurious bias terms, this highlights the potential for erroneous OOD detection when prototype estimates are biased toward spurious features.
For example, in the Waterbird dataset, the background (water or land) represents the spurious feature dimension characterized by distinct mean vectors $\vec{\boldsymbol{\mu}}_{c}^{\text{sp}}$ and $\vec{\boldsymbol{\mu}}_{1-c}^{\text{sp}}$. The difference $\|\vec{\boldsymbol{\mu}}_{c}^{\text{sp}}-\vec{\boldsymbol{\mu}}_{1-c}^{\text{sp}}\|$ may exceed the core bias in OOD distances, causing minority samples with conflicting backgrounds to lie farther from prototypes than some OOD samples.
A.0.7 Why SPROD Mitigates Distance Bias
If prototypes can be accurately estimated to match each group’s distribution, then each ID sample $z_{i}\in\mathcal{S}_{c}^{g}$ can be compared to its corresponding prototype $\hat{\mathbf{p}}(z_{i})=\vec{\mathbf{p}}_{c}^{g}$, where $g\in\{\text{maj},\text{min}\}$.
Because the prototype shares the same spurious basis alignment, the spurious bias is eliminated. The expected squared distance becomes:
$\mathbb{E}\left[\|z_{i}-\hat{\mathbf{p}}(z_{i})\|^{2}\right]=\operatorname{Tr}(\Sigma_{c}^{\text{core}})+\operatorname{Tr}(\Sigma^{\text{irr}}).$
In contrast, the standard prototype-based approach introduces a systematic bias for minority samples:
$\mathbb{E}\left[\|z_{\text{min}}-\vec{\mathbf{p}}_{c}^{\text{maj}}\|^{2}\right]={\color[rgb]{0.5,0.0,0.0}\|\vec{\boldsymbol{\mu}}_{1-c}^{\text{sp}}-\vec{\boldsymbol{\mu}}_{c}^{\text{sp}}\|^{2}}+\operatorname{Tr}(\Sigma_{c}^{\text{core}})+\operatorname{Tr}(\Sigma^{\text{irr}}).$
For an SP-OOD sample, the prototype it is compared against (either group) contains a core component it lacks. Hence the expected distance becomes:
$\mathbb{E}\left[\|z_{\text{OOD}}-\hat{\mathbf{p}}\|^{2}\right]={\color[rgb]{0.0,0.0,0.55}\|\vec{\boldsymbol{\mu}}_{c}^{\text{core}}\|^{2}}+\operatorname{Tr}(\Sigma_{c}^{\text{core}})+\operatorname{Tr}(\Sigma^{\text{irr}}).$
This is strictly greater than the refined ID distance, which includes neither the spurious bias nor the core bias term:
$\mathbb{E}\left[\|z_{\text{OOD}}-\hat{\mathbf{p}}\|^{2}\right]>\mathbb{E}\left[\|z_{i}-\hat{\mathbf{p}}(z_{i})\|^{2}\right].$
Conclusion: SPROD removes the spurious bias term for ID samples and reduces overlap with OOD samples by accounting for their missing core features. This systematic (in expectation) bias elimination leads to tighter ID clusters and more reliable OOD detection.
Appendix B Spurious Correlation in OOD Detection: Literature Review
@@
Outlier exposure based methods include RONF [21], which improves synthetic outlier generation and model fine-tuning using only ID data. It introduces Boundary Feature Mixup to create more realistic virtual outliers by interpolating near decision boundaries, and Optimal Parameter Learning to suppress spurious feature learning during training. At inference, it uses a custom Energy with Energy Discrepancy score to better separate ID from OOD samples without relying on external OOD datasets. Similarly, KIRBY [22] generates hard negative samples by removing class-discriminative regions identified via Class Activation Maps (CAM) and inpainting these regions with background-like content, creating semantically degraded but visually plausible images as surrogate OOD examples. A lightweight rejection network trained on features from both clean and modified images enables strong OOD detection without requiring real OOD data or backbone retraining. Although not directly targeting spurious correlations, ImOOD [23] evaluates robustness under spurious settings, focusing on long-tailed datasets where class imbalance biases OOD detection toward frequent (head) classes. It learns a bias correction term to shift OOD scores per input, improving separation especially for rare (tail) classes.
@@
Post-hoc methods operate after model training to improve OOD detection. Projection Regret (PR) [27] uses partial diffusion and denoising to project inputs onto the ID manifold, measuring semantic novelty by comparing the original image with its projection. A second projection step recursively removes background bias, isolating true semantic differences and enabling object-level OOD detection focused on meaningful changes rather than background similarity. CoVer [28] enhances detection by evaluating model confidence across multiple corrupted versions of the same input (e.g., fog, blur, contrast shifts). Instead of relying on a single prediction, it averages confidence scores over these variants, improving robustness.
@@
Appendix C Comparison with SP-OOD Methods
Most existing OOD detection methods are not directly comparable to SPROD, as they often rely on different assumptions, such as access to auxiliary OOD data, or necessitate retraining the feature backbone with specific objectives. These design choices contrast with SPROD’s post-hoc nature and its objectives of simplicity, efficiency, and broad applicability without model retraining. Nonetheless, to provide a broader context, we include results from a selection of methods that, while not perfectly aligned, have been evaluated under similar SP-OOD conditions. For fairness and consistency, we report the results as published in their original papers, rather than re-implementing them, thereby representing each method according to its best-reported performance. As shown in Table 4, under the most comparable conditions presented, SPROD achieves highly competitive performance, notably without requiring model retraining (when using a ResNet-18 backbone, as applicable to some comparisons) or access to an auxiliary text modality (in the case of methods leveraging CLIP).
@@ Appendix C Comparison with SP-OOD Methods
Appendix D Dataset Details and Examples
Datasets available for studying spurious correlations are generally limited. In the context of OOD detection, this limitation becomes even more pronounced, as the task requires datasets with SP-OOD samples. As a result, the datasets employed must be thoughtfully considered to ensure meaningful evaluation. In this study, our goal is to address a broader range of spurious features present in data, beyond just background features. A distinctive aspect of our experimental design is the inclusion of three additional settings to explore:
@@ Appendix D Dataset Details and Examples
  • CelebA [34]: This dataset is used for a binary classification task with labels $y\in\{\text{blond hair},\text{non-blond hair}\}$, where spurious correlations with gender $\in\{\text{male},\text{female}\}$ are present. It is a real-world dataset, making it suitable for evaluating models in realistic settings. In the dataset, most females have blond hair, and most males have non-blond hair, forming the majority groups, as shown in Figure 11 with colored borders. The minority groups are the opposite combinations. The distribution of groups across classes with varying correlation levels is presented in Table 6. In this setting, SP-OOD samples are those with no hair, but still exhibiting the spurious gender-related features. For our SP-OOD samples, we used bald males, who lack core features but retain spurious (gender) cues.
@@
  • AnimalsMetaCoCo: To create a multi-class, multi-valued spurious attribute setting that reflects realistic and challenging scenarios, we selected a subset of 26 animal classes from the MetaCoCo dataset [71]. The samples were first cleaned through relabeling, removal of duplicates, and deletion of irrelevant images. Subsequently, we defined subattributes to create a total of 8 major concepts for modeling attribute imbalance and spurious features, as illustrated in the first Figure of the paper. We introduce this new dataset as AnimalsMetaCoCo, a refined subset tailored for multi-class, multi-valued spurious feature scenarios. Details of the classes, attributes, and group sample distributions are provided in Table 8. This dataset represents a more realistic scenario, where each class is associated with at least one spurious feature similar to those present in ID data. While the strength of spurious correlations may not be as pronounced as in the other three datasets, AnimalsMetaCoCo introduces several unique and challenging aspects. Specifically, its spurious attributes can take on multiple distinct values, each exhibiting varying degrees of imbalance across classes. Moreover, the class distributions themselves are inherently imbalanced. To construct SP-OOD scenarios, we adopt a leave-2-out strategy: in each round, two classes are treated as SP-OOD, characterized by semantic shifts while retaining at least one spurious attribute shared with ID data, and the remaining classes serve as ID. This setup significantly increases the difficulty of the detection task due to overlapping spurious patterns across environments.
@@
Appendix E Performance on Non-Spurious OOD Datasets